# uniINF: Best-of-Both-Worlds Algorithm for Parameter-Free Heavy-Tailed MABs

### v1

@@ 4.2 Auto-Balancing Learning Rate Scheduling Scheme
The design of the learning rate $S_{t}$ in our algorithm Algorithm 1 is specified in Equation 4 as $S_{t+1}^{2}=S_{t}^{2}+(\ell_{t,i_{t}}^{\text{clip}})^{2}\cdot(1-x_{t,i_{t}})^{2}\cdot(K\log T)^{-1}$. The idea is to balance a Bregman divergence term $\textsc{Div}_{t}$ and a $\Psi$-shifting term $\textsc{Shift}_{t}$ that arise in our regret analysis (roughly corresponding to the terms on the RHS of Equation 5). They allow the following upper bounds as we will include as Lemmas 5 and 9:
@@ 4.3 Adaptive Skipping-Clipping Loss Tuning Technique
For heavy-tailed losses, a unique challenge is that $\mathbb{E}[\ell^{2}_{t,i_{t}}]$, the squared incurred loss appearing in the Bregman divergence term (as shown in Equation 6), does not allow a straightforward upper bound as we only have $\mathbb{E}[\lvert\ell_{t,i_{t}}\rvert^{\alpha}]\leq\sigma^{\alpha}$. A previous workaround is deploying a skipping technique that replaces large loss with 00 (Huang et al.,, 2022). This technique forbids a sudden increase in the divergence term and thus eases the balance between $\textsc{Div}_{t}$ and $\textsc{Shift}_{t}$. However, this skipping is not a free lunch: if we skipped too much, the increase of $S_{t}$ will be slow, which makes the skipping threshold $C_{t,i}$ grow slowly as well — therefore, we end up skipping even more and incurring tremendous skipping error! This eventually stops us from establishing instance-dependent guarantees in stochastic setups.
@@ 5 Main Results
The formal proof of this theorem is provided in Appendix D. As shown in Table 1, our uniINF automatically achieves nearly optimal instance-dependent and instance-independent regret guarantees in stochastic and adversarial environments, respectively. Specifically, under stochastic settings, uniINF achieves the regret upper bound ${\mathcal{O}}\left(K\left(\nicefrac{{\sigma^{\alpha}}}{{\Delta_{\min}}}\right)^{\nicefrac{{1}}{{\alpha-1}}}\log T\cdot\log\nicefrac{{\sigma^{\alpha}}}{{\Delta_{\min}}}\right)$. Compared to the instance-dependent lower bound $\Omega\left(\sum_{i\neq i^{\ast}}(\nicefrac{{\sigma^{\alpha}}}{{\Delta_{i}}})^{\nicefrac{{1}}{{\alpha-1}}}\log T\right)$ given by Bubeck et al., (2013), our result matches the lower bound up to logarithmic factors $\log\nicefrac{{\sigma^{\alpha}}}{{\Delta_{\min}}}$ which is independent of $T$ when all $\Delta_{i}$’s are similar to the dominant sub-optimal gap $\Delta_{\min}$. For adversarial environments, our algorithm achieves an $\widetilde{{\mathcal{O}}}\left(\sigma K^{1-\nicefrac{{1}}{{\alpha}}}T^{\nicefrac{{1}}{{\alpha}}}\right)$ regret, which matches the instance-independent lower bound $\Omega\left(\sigma K^{1-\nicefrac{{1}}{{\alpha}}}T^{\nicefrac{{1}}{{\alpha}}}\right)$ given in Bubeck et al., (2013) up to logarithmic terms. Therefore, the regret guarantees of uniINF are nearly-optimal in both stochastic and adversarial environments .
@@ 5.1 Regret Decomposition
Term I. Benchmark Calibration Error. As in a typical log-barrier analysis (Wei and Luo,, 2018; Ito,, 2021), the Benchmark Calibration Error is not the dominant term. This is because
@@
Term II. Main Regret. By FTRL regret decomposition (see Lemma 29 in the appendix; it is an extension of the classical FTRL bounds (Lattimore and Szepesvári,, 2020, Theorem 28.5)), we have
@@
Notice that the factor $(x_{t,i^{\ast}}-1)$ in the second term is negative and ${\mathcal{F}}_{t-1}$-measurable, and we have
@@
In Sections 5.2, 5.3, and 5.4, we analyze these three key items and introduce our novel analytical techniques for both adversarial and stochastic cases. Specifically, we bound the Bregman divergence terms in Section 5.2 (all formal proofs in Appendix A), the $\Psi$-shifting terms in Section 5.3 (all formal proofs in Appendix B), and the skipping error terms in Section 5.4 (all formal proofs in Appendix C). Afterwards, to get our Theorem 3, putting Theorems 7, 10, and 13 together gives the adversarial guarantee, while the stochastic guarantee follows from a combination of Theorems 8, 11, and 14.
@@ 5.2 Analyzing Bregman Divergence Terms
To simplify notations, we define $w_{t,i}:=\frac{x_{t,i}}{z_{t,i}}-1$, which gives $\textsc{Div}_{t}=S_{t}\sum_{i=1}^{K}(w_{t,i}-\log(w_{t,i}+1))$. Therefore, one natural idea is to utilize the inequality $x-\log(x+1)\leq x^{2}$ for $x\in[-\nicefrac{{1}}{{2}},\nicefrac{{1}}{{2}}]$. To do so, we need to conclude $-\nicefrac{{1}}{{2}}\leq w_{t,i}\leq\nicefrac{{1}}{{2}}$. We develop a novel technical tool to depict ${\bm{z}}_{t}$ and provide an important technical lemma which says ${\bm{z}}_{t}$ is multiplicatively close to ${\bm{x}}_{t}$:
@@
Adversarial Cases.
By definition of $S_{t}$ in Equation 4, we can bound $S_{T+1}$ as in the following lemma.
@@
Stochastic Cases.
For $\operatorname{\mathcal{O}}(\log T)$ bounds, we opt for the first statement of Lemma 5. By definition of $\ell_{t,i_{t}}^{\text{skip}}$, we immediately have $|\ell_{t,i_{t}}^{\text{skip}}|\leq C_{t,i_{t}}=\operatorname{\mathcal{O}}(S_{t}(1-x_{t,i_{t}})^{-1})$. Further using $\mathbb{E}[\lvert\ell_{t,i_{t}}\rvert^{\alpha}]\leq\sigma^{\alpha}$, we have $\mathbb{E}[\textsc{Div}_{t}\mid\mathcal{F}_{t-1}]={\mathcal{O}}(S_{t}^{1-\alpha}\sigma^{\alpha}(1-x_{t,i^{\ast}}))$ (formalized as Lemma 20 in the appendix).
We can now perform a stopping-time argument for the sum of Bregman divergence terms in stochastic cases. Briefly, we pick a fixed constant $M>0$. The expected sum of $\textsc{Div}_{t}$’s on those $\{t\mid S_{t}\geq M\}$ is then within $\operatorname{\mathcal{O}}(M^{1-\alpha}\cdot T)$ according to Lemma 5. On the other hand, we claim that the sum of $\textsc{Div}_{t}$’s on those $\{t\mid S_{t}<M\}$ can also be well controlled because $\mathbb{E}[\textsc{Div}_{t}\mid\mathcal{F}_{t-1}]={\mathcal{O}}(\mathbb{E}[S_{t}^{1-\alpha}\sigma^{\alpha}(1-x_{t,i^{\ast}})\mid{\mathcal{F}}_{t-1}])$. The detailed analysis is presented in Section A.6, and we summarize it as follows. Therefore, the sum of Bregman divergences in stochastic environments is also well-controlled.
@@ 5.3 Analyzing $\Psi$-Shifting Terms
Adversarial Cases.
Again, we utilize the $\mathbb{E}[S_{T+1}]$ bound in Lemma 6 and get the following theorem.
@@
Stochastic Cases.
Still similar to $\textsc{Div}_{t}$’s, we condition on ${\mathcal{F}}_{t-1}$ and utilize $\lvert\ell_{t,i_{t}}^{{\text{clip}}}\rvert\leq C_{t,i_{t}}$ to get $\mathbb{E}\left[\textsc{Shift}_{t}\mid{\mathcal{F}}_{t-1}\right]={\mathcal{O}}(S_{t}^{1-\alpha}\sigma^{\alpha}(1-x_{t,i^{\ast}}))$ (formalized as Lemma 24 in the appendix). Thus, for stochastic environments, a stopping-time argument similar to that of $\textsc{Div}_{t}$ yields Theorem 11.
@@ 5.4 Analyzing Sub-Optimal Skipping Losses
For skipping errors, we need a more dedicated stopping-time analysis in both adversarial and stochastic cases. Different from previous sections, it is now non-trivial to bound the sum of $\textsc{SkipErr}_{t}$’s on those $\{t\mid S_{t}<M\}$ where $M$ is the stopping-time threshold. However, a key observation is that whenever we encounter a non-zero $\textsc{SkipErr}_{t}$, $S_{t+1}$ will be $\Omega(1)$-times larger than $S_{t}$, thanks to the adaptive skipping-clipping technique. Thus the number of non-zero $\textsc{SkipErr}_{t}$’s before $S_{t}$ reaching $M$ is small. Equipped with this observation, we derive the following lemma, whose formal proof is included in Section C.1.
@@
Equipped with this novel stopping-time analysis, it only remains to pick a proper threshold $M$ for Lemma 12. It turns out that for adversarial and stochastic cases, we have to pick different $M$’s. Specifically, we achieve the following two theorems, whose proofs are in Sections C.2 and C.3.
@@
$\mathbb{E}\left[\sum_{t=1}^{T}|\textsc{SkipErr}_{t}|\cdot\mathbbm{1}[i_{t}\neq i^{\ast}]\right]={\mathcal{O}}\left(K\log T\cdot\sigma^{\frac{\alpha}{\alpha-1}}\Delta_{\min}^{-\frac{1}{\alpha-1}}\cdot\log(\sigma^{\alpha}/\Delta_{\min})+\frac{{\mathcal{R}}_{T}}{4}\right).$
@@ 6 Conclusion
Appendix A Bregman Divergence Terms: Omitted Proofs in Section 5.2

### v2

@@ 4.2 Auto-Balancing Learning Rate Scheduling Scheme
We design the our auto-balancing learning rate $S_{t}$ in Equation 4. The idea is to balance a Bregman divergence term $\textsc{Div}_{t}$ and a $\Psi$-shifting term $\textsc{Shift}_{t}$ that arise in our regret analysis (roughly corresponding to the terms on the RHS of Equation 5). They allow the following upper bounds as we will include as Lemmas 5 and 9:
@@ 4.3 Adaptive Skipping-Clipping Loss Tuning Technique
For heavy-tailed losses, a unique challenge is that $\mathbb{E}[\ell^{2}_{t,i_{t}}]$, the squared incurred loss appearing in the Bregman divergence term (as shown in Equation 6), does not allow a straightforward upper bound as we only have $\mathbb{E}[\lvert\ell_{t,i_{t}}\rvert^{\alpha}]\leq\sigma^{\alpha}$. A previous workaround is deploying a skipping technique that replaces large loss with 00 (Huang et al., 2022). This technique forbids a sudden increase in the divergence term and thus eases the balance between $\textsc{Div}_{t}$ and $\textsc{Shift}_{t}$. However, this skipping is not a free lunch: if we skipped too much, the increase of $S_{t}$ will be slow, which makes the skipping threshold $C_{t,i}$ grow slowly as well — therefore, we end up skipping even more and incurring tremendous skipping error! This eventually stops us from establishing instance-dependent guarantees in stochastic setups.
@@ 5 Main Results
The formal proof of this theorem is provided in Appendix F. As shown in Table 1, our uniINF automatically achieves nearly optimal instance-dependent and instance-independent regret guarantees in stochastic and adversarial environments, respectively. Specifically, under stochastic settings, uniINF achieves the regret upper bound ${\mathcal{O}}\left(K\left(\nicefrac{{\sigma^{\alpha}}}{{\Delta_{\min}}}\right)^{\nicefrac{{1}}{{\alpha-1}}}\log T\cdot\log\nicefrac{{\sigma^{\alpha}}}{{\Delta_{\min}}}\right)$. Compared to the instance-dependent lower bound $\Omega\left(\sum_{i\neq i^{\ast}}(\nicefrac{{\sigma^{\alpha}}}{{\Delta_{i}}})^{\nicefrac{{1}}{{\alpha-1}}}\log T\right)$ given by Bubeck et al. (2013), our result matches the lower bound up to logarithmic factors $\log\nicefrac{{\sigma^{\alpha}}}{{\Delta_{\min}}}$ which is independent of $T$ when all $\Delta_{i}$’s are similar to the dominant sub-optimal gap $\Delta_{\min}$. For adversarial environments, our algorithm achieves an $\widetilde{{\mathcal{O}}}\left(\sigma K^{1-\nicefrac{{1}}{{\alpha}}}T^{\nicefrac{{1}}{{\alpha}}}\right)$ regret, which matches the instance-independent lower bound $\Omega\left(\sigma K^{1-\nicefrac{{1}}{{\alpha}}}T^{\nicefrac{{1}}{{\alpha}}}\right)$ given in Bubeck et al. (2013) up to logarithmic terms. Therefore, the regret guarantees of uniINF are nearly-optimal in both stochastic and adversarial environments .
@@ 5.1 Regret Decomposition
Term I. Benchmark Calibration Error. As in a typical log-barrier analysis (Wei & Luo, 2018; Ito, 2021b), the Benchmark Calibration Error is not the dominant term. This is because
@@
Term II. Main Regret. By FTRL regret decomposition (see Lemma 29 in the appendix; it is an extension of the classical FTRL bounds (Lattimore & Szepesvári, 2020, Theorem 28.5)), we have
@@
To summarize, the regret ${\mathcal{R}}_{T}$ decomposes into the sum of Bregman divergence terms $\mathbb{E}[\textsc{Div}_{t}]$, $\Psi$-shifting terms $\mathbb{E}[\textsc{Shift}_{t}]$, and sub-optimal skipping losses $\mathbb{E}[|\textsc{SkipErr}_{t}|\cdot\mathbbm{1}[i_{t}\neq i^{\ast}]]$, namely
@@
In Sections 5.2, 5.3, and 5.4, we analyze these three key items and introduce our novel analytical techniques for both adversarial and stochastic cases. Specifically, we bound the Bregman divergence terms in Section 5.2 (all formal proofs in Appendix C), the $\Psi$-shifting terms in Section 5.3 (all formal proofs in Appendix D), and the skipping error terms in Section 5.4 (all formal proofs in Appendix E). Afterwards, to get our Theorem 3, putting Theorems 7, 10, and 13 together gives the adversarial guarantee, while the stochastic guarantee follows from a combination of Theorems 8, 11, and 14.
@@ 5.2 Analyzing Bregman Divergence Terms
To simplify notations, we define $w_{t,i}:=\frac{x_{t,i}}{z_{t,i}}-1$, which gives $\textsc{Div}_{t}=S_{t}\sum_{i=1}^{K}(w_{t,i}-\log(w_{t,i}+1))$. Therefore, one natural idea is to utilize the inequality $x-\log(x+1)\leq x^{2}$ for $x\in[-\nicefrac{{1}}{{2}},\nicefrac{{1}}{{2}}]$. To do so, we need to ensure $-\nicefrac{{1}}{{2}}\leq w_{t,i}\leq\nicefrac{{1}}{{2}}$, i.e., ${\bm{z}}_{t}$ is multiplicatively close to ${\bm{x}}_{t}$. If the losses are bounded, this property is achieved in previously papers (Lee et al., 2020a; Jin & Luo, 2020). However, it is non-trivial for heavy-tailed (thus unbounded) losses. By refined analysis on the calculation of ${\bm{x}}$ and ${\bm{z}}$ by KKT conditions, we develop novel technical tools to depict ${\bm{z}}_{t}$ and provide the following important lemma in HTMABs. We refer readers to the proof of Lemma 15 for more technical details.
@@
Adversarial Cases. By definition of $S_{t}$ in Equation 4, we can bound $S_{T+1}$ as in the following lemma.
@@
Stochastic Cases. For $\operatorname{\mathcal{O}}(\log T)$ bounds, we opt for the first statement of Lemma 5. By definition of $\ell_{t,i_{t}}^{\text{skip}}$, we immediately have $|\ell_{t,i_{t}}^{\text{skip}}|\leq C_{t,i_{t}}=\operatorname{\mathcal{O}}(S_{t}(1-x_{t,i_{t}})^{-1})$. Further using $\mathbb{E}[\lvert\ell_{t,i_{t}}\rvert^{\alpha}]\leq\sigma^{\alpha}$, we have $\mathbb{E}[\textsc{Div}_{t}\mid\mathcal{F}_{t-1}]={\mathcal{O}}(S_{t}^{1-\alpha}\sigma^{\alpha}(1-x_{t,i^{\ast}}))$ (formalized as Lemma 20 in the appendix).
We can now perform a stopping-time argument for the sum of Bregman divergence terms in stochastic cases. Briefly, we pick a fixed constant $M>0$. The expected sum of $\textsc{Div}_{t}$’s on those $\{t\mid S_{t}\geq M\}$ is then within $\operatorname{\mathcal{O}}(M^{1-\alpha}\cdot T)$ according to Lemma 5. On the other hand, we claim that the sum of $\textsc{Div}_{t}$’s on those $\{t\mid S_{t}<M\}$ can also be well controlled because $\mathbb{E}[\textsc{Div}_{t}\mid\mathcal{F}_{t-1}]={\mathcal{O}}(\mathbb{E}[S_{t}^{1-\alpha}\sigma^{\alpha}(1-x_{t,i^{\ast}})\mid{\mathcal{F}}_{t-1}])$. The detailed analysis is presented in Section C.6, and we summarize it as follows. Therefore, the sum of Bregman divergences in stochastic environments is also well-controlled.
@@ 5.3 Analyzing $\Psi$-Shifting Terms
Adversarial Cases. Again, we utilize the $\mathbb{E}[S_{T+1}]$ bound in Lemma 6 and get the following theorem.
@@
Stochastic Cases. Still similar to $\textsc{Div}_{t}$’s, we condition on ${\mathcal{F}}_{t-1}$ and utilize $\lvert\ell_{t,i_{t}}^{{\text{clip}}}\rvert\leq C_{t,i_{t}}$ to get $\mathbb{E}\left[\textsc{Shift}_{t}\mid{\mathcal{F}}_{t-1}\right]={\mathcal{O}}(S_{t}^{1-\alpha}\sigma^{\alpha}(1-x_{t,i^{\ast}}))$ (formalized as Lemma 24 in the appendix). Thus, for stochastic environments, a stopping-time argument similar to that of $\textsc{Div}_{t}$ yields Theorem 11.
@@ 5.4 Analyzing Sub-Optimal Skipping Losses
For skipping errors, we need a more dedicated stopping-time analysis in both adversarial and stochastic cases. Different from previous sections, it is now non-trivial to bound the sum of $\textsc{SkipErr}_{t}$’s on those $\{t\mid S_{t}<M\}$ where $M$ is the stopping-time threshold. However, a key observation is that whenever we encounter a non-zero $\textsc{SkipErr}_{t}$, $S_{t+1}$ will be $\Omega(1)$-times larger than $S_{t}$. Thus the number of non-zero $\textsc{SkipErr}_{t}$’s before $S_{t}$ reaching $M$ is small. Equipped with this observation, we derive the following lemma, whose formal proof is included in Section E.1.
@@
Equipped with this novel stopping-time analysis, it only remains to pick a proper threshold $M$ for Lemma 12. It turns out that for adversarial and stochastic cases, we have to pick different $M$’s. Specifically, we achieve the following two theorems, whose proofs are in Sections E.2 and E.3.
@@
$\mathbb{E}\left[\sum_{t=1}^{T}|\textsc{SkipErr}_{t}|\cdot\mathbbm{1}[i_{t}\neq i^{\ast}]\right]={\mathcal{O}}\left(K\log T\cdot\sigma^{\frac{\alpha}{\alpha-1}}\Delta_{\min}^{-\frac{1}{\alpha-1}}\cdot\log\left(\frac{\sigma^{\alpha}}{\Delta_{\min}}\right)+\frac{{\mathcal{R}}_{T}}{4}\right).$
@@ 6 Conclusion
Reproducibility Statement
This paper is primarily theoretical in nature, with the main contributions being the development of uniINF, the Best-of-Both-Worlds algorithm for Parameter-Free Heavy-Tailed Multi-Armed Bandits, and the introduction of several innovative algorithmic components and analytical tools. To ensure reproducibility, we have provided clear and detailed explanations of our methods, assumptions, and the development process of the algorithm throughout the paper. Complete proofs of our claims and results are included in the appendices. We have made every effort to present our research in a transparent and comprehensive manner to enable other researchers to understand and replicate our results.
Acknowledgments
We thank the anonymous reviewers for their detailed reviews, from which we benefit greatly. This work was supported by the National Natural Science Foundation of China Grants 52450016 and 52494974, and in part by the National Science Foundation of China under Grant 623B1015.
Appendix A Additional Related Works
In this section, we present a detailed version of literature review to include more discussion and related works.
A.1 Multi-Armed Bandits
The Multi-Armed Bandits (MAB) problem (Thompson, 1933; Slivkins et al., 2019; Lattimore & Szepesvári, 2020; Bubeck et al., 2012) is an important theoretical framework for addressing the exploration-exploitation trade-off inherent in online learning. Various studies discovered different setups for the loss (or reward) signals. For example, most works focus on standard bounded losses where the loss of each arm belongs to a bounded range such as $[0,1]$ (Auer et al., 2002; Wei & Luo, 2018; Lee et al., 2020a; Jin et al., 2023). Srinivas et al. (2009); Wu et al. (2015); Deng et al. (2022) assumed the losses following Gaussian or sub-Gaussian distribution. Another branch of works called scale-free MABs consider the case when the losses’ range is unknown to the agent (Putta & Agrawal, 2022; Chen & Zhang, 2023; 2024; Hadiji & Stoltz, 2023). In this paper, we consider the heavy-tailed loss signals where each arm’s loss do not allow bounded variances but instead have their $\alpha$-th moment bounded by some constant $\sigma^{\alpha}$, which is called the Heavy-Tailed Multi-Armed Bandits (HTMABs).
A.2 Heavy-Tailed MABs
HTMABs were introduced by Bubeck et al. (2013), who gave both instance-dependent and instance-independent lower and upper bounds under stochastic assumptions. Various efforts have been devoted in this area since then. To exemplify, Wei & Srivastava (2020) removed a sub-optimal $(\log T)^{1-1/\alpha}$ factor in the instance-independent upper bound; Yu et al. (2018) developed a pure exploration algorithm for HTMABs; Medina & Yang (2016), Kang & Kim (2023), and Xue et al. (2024) considered the linear HTMAB problem; and Dorn et al. (2024) specifically investigated the case where the heavy-tailed reward distributions are presumed to be symmetric. Nevertheless, all these works focused on stochastic environments and required the prior knowledge of heavy-tail parameters $(\alpha,\sigma)$. In contrast, this paper focuses on both-of-both-worlds algorithms in parameter-free HTMABs, which means i) the loss distributions can possibly be non-stationary, and ii) the true heavy-tail parameters $(\alpha,\sigma)$ remain unknown.
A.3 Best-of-Both-Worlds Algorithms
Bubeck & Slivkins (2012) pioneered the study of Best-of-Both-Worlds bandit algorithms, and was followed by various improvements including EXP3-based approaches (Seldin & Slivkins, 2014; Seldin & Lugosi, 2017), FTPL-based approaches (Honda et al., 2023; Lee et al., 2024), and FTRL-based approaches (Wei & Luo, 2018; Zimmert & Seldin, 2019; Jin et al., 2023). Equipped with data-dependent learning rates, Ito (2021a; b); Tsuchiya et al. (2023); Ito & Takemura (2023); Kong et al. (2023) achieve Best-of-Both-Worlds algorithms in various online learning problems. When the loss distribution can be heavy-tailed, Huang et al. (2022) gave an algorithm HTINF that achieves Best-of-Both-Worlds property under the known-$(\alpha,\sigma)$ assumption. Unfortunately, without access to these true parameters, their alternative algorithm OptHTINF failed to achieve near-optimal regret guarantees in either adversarial or stochastic environments.
A.4 Parameter-Free HTMABs
Another line of research aimed at getting rid of the prior knowledge of $\alpha$ or $\sigma$, which we call Parameter-Free HTMABs. Along this line, Kagrecha et al. (2019) presented the GSR method to identify the optimal arm in HTMAB without any prior knowledge. In terms of regret minimization, Lee et al. (2020b) and Lee & Lim (2022) considered the case when $\sigma$ is unknown. Genalti et al. (2024) were the first to achieve the parameter-free property while maintaining near-optimal regret. However, all these algorithms fail in adversarial environments – not to mention the best-of-both-worlds property which requires optimality in both stochastic and adversarial environments.
Appendix B Discussions on Assumption 1 (Truncated Non-Negative Loss)
In this section, we detailedly discuss the existing results on Assumption 1 in HTMABs works. It is important to note that the truncated non-negativity loss assumption is not unique to our study. In fact, it was also used in Huang et al. (2022) and Genalti et al. (2024), and is indeed a relaxation of the more common "non-negative losses" assumption found in the MAB literature (Auer et al., 2002; Lee et al., 2020a; Jin et al., 2023). For a better comparison, we summarize the exisiting results in HTMABs for the relationship between the prior knowledge of heavy-tailed parameters $(\sigma,\alpha)$ and Assumption 1 in the following table.
Table 2: Impact of Assumption 1 on BoBW Guarantees for HTMABs
Assumptions Known $(\sigma,\alpha)$ Unknown $(\sigma,\alpha)$
With Assumption 1 HTINF achieves BoBW uniINF achieves BoBW
(Huang et al., 2022) (Our Work)
Weaker than Assumption 1? Inferred by Cheng et al. (2024) Open Problem
(discussed below)
Without Any Assumption SAO-HT achieves BoBW No BoBW possible
(Cheng et al., 2024) (Genalti et al., 2024, Theorems 2 & 3)
The truncated non-negative loss assumption is first introduced by Huang et al. (2022), and their algorithm HTINF achieves nearly optimal automatically in both stochasitc and adversarial environments with Assumption 1 and prior knowledge of heavy-tailed parameters $(\sigma,\alpha)$, as shown in Table 2.
Recently, Cheng et al. (2024) justified that when parameters $(\sigma,\alpha)$ are known, BoBW is achievable without any assumptions (that is, Assumption 1 is redundant when parameters are known). Specifically, the algorithm SAO-HT presented by Cheng et al. (2024) achieves nearly optimal in adversarial and stochastic cases up to some extra logarithm factors. This result implies that Assumption 1 is redundant for optiaml HTMABs when parameters are known.
When it comes to parameter-free setups, Theorems 2 and 3 by Genalti et al. (2024) highlight that achieving optimal worst-case regret guarantees is impossible without any assumptions. Moreover, Genalti et al. (2024) present an efficient algorithm Adar-UCB which achieve nearly minimax optimal on both instance-dependent and instance-independent regret bounds, under the stochastic environments and Assumption 1.
Therefore, we adopt Assumption 1 when developing our BoBW algorithms for parameter-free HTMABs. Our algorithm, uniINF (Algorithm 1), achieves nearly optimal automatically in both stochastic and adversarial settings without knowing the parameters $(\sigma,\alpha)$ a-priori. Our result also demonstrates that Assumption 1 is sufficient for BoBW guarantees in parameter-free HTMABs.
As summarized by Table 2, while Assumption 1 has been validated as sufficient for achieving BoBW guarantees in both known and unknown parameter settings, the exploration for weaker assumptions continues to be an intersting topic for future research.
Appendix C Bregman Divergence Terms: Omitted Proofs in Section 5.2